# A Probabilistic Parking Process and Labeled IDLA

### v1

@@ 4. Probabilistic Parking with Open Boundaries
Figure 3. The color of each cell $(p,y)$ in the image above corresponds to the number of parking functions $\alpha\in\mathrm{PF}_{n}(\mathrm{id})$ where the probability that $\alpha$ parks with parameter $p$ is at most $y$, where $y$ is calculated according to Theorem 4.1. For example, when $p=0$, the probability that $\alpha$ parks is always 00 and the pixels are dark red for all $y$.
@@
Solving this system of equations when $p\neq\frac{1}{2}$ (for the technique used refer to; [25, Chapter 12.2]), assuming that car $i$ starts in position $s\in[n]$ where $1\leq s\leq i$, we have that
@@ 4.1. Expected Time to Park
denote the random stopping time which is the first time that car $i$ hits position $i$ or position 00. Recall that we let $\tau_{\alpha}$ denote the time it takes for all cars to park. In other words, $\tau_{\alpha}$ is the stopping time such that
@@
Recall that, by the dynamics of the process, in order for all of the cars to successfully park we must have that each car $i$ parks on its “first attempt”. In other words, if a car encounters either position $n+1$ or position 00 before position $i$, then the entire parking process was unsuccessful. Thus, we have the following theorem giving the expected time for a single car to park, given that all previous cars park.
@@
Let $g^{i}_{s}\coloneqq\mathbb{E}_{s}\left[\tau_{\alpha}^{i}\mid\prod_{j=1}^{i}X_{\alpha}^{j}=1\right]$ where we use $\mathbb{E}_{s}$ to denote the expectation of the process starting from $s$. In other words $g^{i}_{s}$ is the expected time it takes for car $i$ to successfully park when starting from position $s$, and thus, $g^{i}_{0}=g^{i}_{i}=0$. Furthermore, letting $\gamma$ be a path from $s$ to $i$, we have that
@@
$\sum_{\ell=1}^{\infty}\sum_{\begin{subarray}{c}\begin{subarray}{c}\begin{subarray}{c}\begin{subarray}{c}|\gamma|=\ell\\ \gamma_{0}=s\end{subarray}\\ \gamma_{1}=s+1\end{subarray}\\ \vdots\end{subarray}\\ \gamma_{\ell}=i\end{subarray}}^{\infty}(\ell-1)\Pr\left[\gamma\cap\left\{Z_{\tau_{\alpha}^{i}}^{i}=i,Z_{0}^{i}=s\right\}\right]+\Pr\left[\gamma\cap\left\{Z_{\tau_{\alpha}^{i}}^{i}=i,Z_{0}^{i}=s\right\}\right]$
$=\sum_{\ell=1}^{\infty}p\sum_{\begin{subarray}{c}\begin{subarray}{c}\begin{subarray}{c}|\beta|=\ell-1\\ \beta=s+1\end{subarray}\\ \vdots\end{subarray}\\ \beta_{\ell}=i\end{subarray}}^{\infty}(\ell-1)\Pr\left[\beta\cap\left\{Z_{\tau_{\alpha}^{i}}^{i}=i,Z_{0}^{i}=s+1\right\}\right]+\Pr\left[\beta\cap\left\{Z_{\tau_{\alpha}^{i}}^{i}=i,Z_{0}^{i}=s+1\right\}\right]$
[…]
$\hskip 36.135pt+\Pr\left[\beta\big{|}\left\{Z_{\tau_{\alpha}^{i}}^{i}=i,Z_{0}^{i}=s+1\right\}\right]\Pr\left[\left\{Z_{\tau_{\alpha}^{i}}^{i}=i,Z_{0}^{i}=s+1\right\}\right]$
where the equality in ($\ast$) uses the definition of conditional expectation. By the definition of $w^{i}_{s}$ and $g^{i}_{s}$ this last line is just
$\begin{split}&=\sum_{\ell=1}^{\infty}pw^{i}_{s+1}\sum_{\begin{subarray}{c}\begin{subarray}{c}\begin{subarray}{c}|\beta|=\ell-1\\ \beta_{1}=s+1\end{subarray}\\ \vdots\end{subarray}\\ \beta_{\ell}=i\end{subarray}}^{\infty}(\ell-1)\Pr\left[\beta\big{|}\left\{Z_{\tau_{\alpha}^{i}}^{i}=i,Z_{0}^{i}=s+1\right\}\right]+\Pr[\beta\big{|}\left\{Z_{\tau_{\alpha}^{i}}^{i}=i,Z_{0}^{i}=s+1\right\}]\\ &=pw^{i}_{s+1}\left(g_{s+1}^{i}+1\right)\,.\end{split}$
@@
(4.4) $g_{s}^{i}w_{s}=pw_{s+1}\left(g_{s+1}^{i}\right)+qw_{s-1}\left(g_{s-1}^{i}\right)+w_{s}\,.$
Solving the system of linear equations when $p\neq q$ yields
@@
The set of parking functions that are weakly increasing are in bijection with those that are weakly decreasing via the involution that takes $i\mapsto n-i+1$, with the added condition that we reverse the order of entrance into the street. That is, the cars would now enter from the right, and perform the parking protocol moving toward spot 1. This involution and change of parking direction in the parking protocol also give a bijection between the set of parking functions whose outcome is the identity permutation and the set of (right to left) parking functions whose outcome is the longest word, namely the permutation $n(n-1)(n-2)\cdots 321$ written in one-line notation. We denote the set of all such parking functions by $\operatorname{PF}_{n}(n\cdots 321)$ and note that they are characterized as the set of parking functions $\beta\in[n]^{n}$ such that the $i$th entry $\beta_{i}$ is at least $n-i+1$ for all $i\in[n]$. By way of this bijection, all of our previous arguments and results for the probabilistic parking protocols (both in the open boundary and the unbounded models) work analogously when starting with the set $\operatorname{PF}_{n}(n\cdots 321)$, of parking functions with outcome $n\cdots 321$.
@@ 4.2. Combinatorial Derivation of Expected Time to Park
First, we consider the number of paths from spot $1$ to spot $i$ in $t$ steps. Following [36], we let $a_{j,k}$ be the number of paths to spot $k+1$ using $j$ steps to the left. We can think of $k$ as the net number of spots that the car has moved to the right, so $k+2j$ is the total number of steps the paths enumerated by $a_{j,k}$ requires. There is one path to any spot using only right steps, so $a_{0,k}=1$. We also have that, analogously to (3.1), the number of paths can be determined by the recurrence relation $a_{j,k}=a_{j,k-1}+a_{j-1,k+1}$. This is because a path to spot $s$ can be obtained from either a path landing at spot $s-1$ followed by one right step, or a path landing in spot $s+1$ followed by one left step. Unlike in (3.1) however, a path in the open boundary case cannot end in spot 00. Lastly, since car $i$ always parks in spot $i$ before reaching spot $i+1$, we have that the number of paths landing in spot $k+1$ in $k+2j$ steps is the sequence $a_{j,k}$ from [36] given by
(4.5) $a_{i,j}=\begin{cases}1&\mbox{ if }j=0\\ a_{j-1,1}&\mbox{ if }k=0,\ 1\leq j\\ a_{j,k-1}+a_{j-1,k+1}&\text{ if }1\leq k\leq i,\ 1\leq j\\ a_{j,\ell}&\mbox{ if }\ell=i+1,\ 1\leq j.\end{cases}$
For any $i$, László Németh and László Szalay refer to the sequence $(a_{j,k})_{j=0}^{\infty}$ as $A_{k}^{(i)}$. Using this notation, the number of paths that car $i$ has starting at spot 1, moving between spot $1$ and $i-1$ in the first $k+2j-1$ steps of the walk and ending at spot $i$ at step $k+2j$ step is the $j$th element in the sequence $A_{i+2}^{(i+2)}$. We use this to find the expected time it takes for car $i$ to travel from spot 1 to spot $i$, given that car $i$ eventually does park which we again denote by $g_{1}^{i}$. We remark that the approach in this section is very similar to that used by McMullen [34], however in their case the terms of the sequence $A_{k}^{(i)}$ are Catalan convolutions because of the different boundary condition. Using the recurrence relation in (4.5), we have the following theorem.
@@
We have remarked that the $j$th element of $A_{i-3}^{(i-3)}$ is the number of paths from 1 to $i$ with $j$ steps to the left. Since we know any path from $1$ to $i$ has $i-1$ more right steps than left steps, we have $j+j+i-1$ total steps in the walk. The probability of such a path occurring from the set of any path is $q^{j}$ times $p^{j+i-1}$ for the left and right steps, respectively. Where we again use $w_{1}^{i}$ to denote the probability of car $i$ parking in spot $i$ before leaving the segment. Thus, since we only want to sample from the functions that park we multiply $p^{k+i-1}q^{k}$ by $1/w_{1}^{i}$ for the probability of selecting a path with $j$ left steps and $12j+i-$ 1 total steps, from those which never go to the left of spot 1. ∎
@@
Given a parking function $\alpha=(\alpha_{1},\alpha_{2},\ldots,\alpha_{n})\in\mathrm{PF}_{n}(\mathrm{id})$, the expected time for car $i$ to park in spot $i$ starting from spot $i-1$ provided it parks and all the previous cars have parked under the probabilistic parking protocol with open boundaries is given by
@@ 5. The Parking Process is Negatively Correlated
In this section, we prove that the indicator random variables $X^{i}_{\alpha}$, which are equal to one if car $i$ parks during the probabilistic parking protocol with open boundaries starting with preference list $\alpha$, and zero otherwise, are pairwise negatively correlated.
@@
In the case of our parking protocol with open boundaries, this is exactly as one might expect. Once a car parks, there are fewer spots in the parking lot and so the probability a car is able to park before leaving the segment should only remain the same or decrease. Note that in the context of the unbounded model on the other hand, the analogously defined indicator variables $X^{i}_{\alpha}$ are not negatively correlated. This is because, a car parks under the unbounded model starting with $\alpha\in\mathrm{PF}_{n}(\mathrm{id})$ if and only if it is able to hit position $i$ before drifting off to infinity. However in the case that car $i$ drifts off to infinity, none of the cars in the set $[i+1,n]$ are able to begin the parking process, and thus park with probability zero. We will also give a proof of this fact at the end of the section.
@@
Consider two indices $i,j\in[n]$ with $i<j$. First notice that the probability of car $i$ parking is not influenced by car $j$ whenever $i<j$. Thus the left hand side of (5.1) gives
(5.2) $\Pr[X^{i}_{\alpha}X^{j}_{\alpha}=1]=\Pr[X^{i}_{\alpha}=1]\Pr[X^{j}_{\alpha}=1|X^{i}_{\alpha}=1]\,.$
Therefore, what is left to show is that
$\begin{split}\Pr[X^{i}_{\alpha}=1]\Pr[X^{j}_{\alpha}=1|X^{i}_{\alpha}=1]&\stackrel{{\scriptstyle?}}{{\leq}}\Pr[X^{i}_{\alpha}=1]\Pr[X^{j}_{\alpha}=1]\,.\\ \end{split}$
By Bayes’ rule, we have
$\Pr[X^{i}_{\alpha}=1]\Pr[X^{j}_{\alpha}=1|X^{i}_{\alpha}=1]=\Pr[X^{j}_{\alpha}=1]\Pr[X^{i}_{\alpha}=1|X^{j}_{\alpha}=1]$
but $\Pr[X^{i}_{\alpha}=1|X^{j}_{\alpha}=1]=\Pr[X^{i}_{\alpha}=1]$ because the probability of an earlier car parking is independent of that of a later car. Thus, we have
$\Pr[X^{i}_{\alpha}X^{j}_{\alpha}=1]=\Pr[X^{i}_{\alpha}=1]\Pr[X^{j}_{\alpha}=1|X^{i}_{\alpha}=1]=\Pr[X^{i}_{\alpha}=1]\Pr[X^{j}_{\alpha}=1]\,,$
concluding the proof. ∎
@@
$\prod\limits_{m=i_{1}}^{i_{s}}\Pr\left[X^{m}_{\alpha}=1\mid\prod\limits_{j=i_{1}}^{m-1}X^{j}_{\alpha}=1\right]\stackrel{{\scriptstyle?}}{{\leq}}\prod\limits_{i\in S}\Pr[X^{i}_{\alpha}=1]$
as the left hand side of the above inequality is precisely equal to $\Pr\left[\prod\limits_{i\in S}X^{i}_{\alpha}=1\right]$ by the definition of conditional probability and the probabilistic parking protocol. Thus, we proceed by induction on the size of the set $S$. For the base case, when $S$ is of size 1, the desired inequality holds directly. In addition, for $S$ containing two cars, Theorem 5.3 proves the inequality holds. Now assume for induction that the result holds for $|S|=k$. We now show the inequality holds for $|S|=k+1$. Again, let $\bar{S}\coloneqq\{i_{1},\dots,i_{k+1}\}$ where $i_{1}<\dots<i_{k+1}$. Using Bayes’ rule, we have
$\begin{split}\prod\limits_{m=i_{1}}^{i_{k+1}}\Pr\left[X^{m}_{\alpha}=1\mid\prod\limits_{j=i_{1}}^{m-1}X^{j}_{\alpha}=1\right]&=\Pr\left[\prod\limits_{m=i_{1}}^{i_{k}}X^{m}_{\alpha}=1\right]\Pr\left[X^{i_{k+1}}_{\alpha}=1\mid\prod\limits_{j=i_{1}}^{i_{k}}X^{j}_{\alpha}=1\right]\\ &=\Pr\left[X^{i_{k+1}}_{\alpha}=1\right]\Pr\left[\prod\limits_{j=i_{1}}^{i_{k}}X^{j}_{\alpha}=1\mid X^{i_{k+1}}_{\alpha}=1\right]\\ &=\Pr\left[X^{i_{k+1}}_{\alpha}=1\right]\Pr\left[\prod\limits_{j=i_{1}}^{k+1}X^{j}_{\alpha}=1\right]\,.\\ \end{split}$
Thus, using the induction hypothesis, we have
$\begin{split}\prod\limits_{m=i_{1}}^{i_{k+1}}\Pr\left[X^{m}_{\alpha}=1\mid\prod\limits_{j=i_{1}}^{m-1}X^{j}_{\alpha}=1\right]&\leq\Pr\left[X^{i_{k+1}}_{\alpha}=1\right]\prod\limits_{j=i_{1}}^{i_{k}}\Pr\left[X^{j}_{\alpha}=1\right]\\ \end{split}$
yielding the desired result. ∎
@@
By our assumption, $i\neq 1$. Furthermore, by Theorem 3.1 we know that the probability that any particular car (which is not the first car) parks with probability 1, depends on the parameter $p$. Thus, it is not true that this inequality always holds, and so the probabilistic parking process with the unbounded boundary condition, is not negatively correlated in general. ∎

### v2

@@ 4. Probabilistic Parking with Open Boundaries
Figure 4. The color of each cell $(p,y)$ in the image above corresponds to the number of parking functions $\alpha\in\mathrm{PF}_{n}(\mathrm{id})$ where the probability that $\alpha$ parks with parameter $p$ is at most $y$, where $y$ is calculated according to Theorem 4.1. For example, when $p=0$, the probability that $\alpha$ parks is always 0 and the pixels are dark red for all $y$.
@@
Solving this system of equations when $p\neq\frac{1}{2}$ (for the technique used refer to: [grinstead2006grinstead, Chapter 12.2]), assuming that car $i$ starts in position $s\in[n]$ where $1\leq s\leq i$, we have that
@@ 4.1. Expected Time to Park
denote the random stopping time which is the first time that car $i$ hits position $i$ or position 0. Recall that we let $\tau_{\alpha}$ denote the time it takes for all cars to park. In other words, $\tau_{\alpha}$ is the stopping time such that
@@
Recall that, by the dynamics of the process, in order for all of the cars to successfully park we must have that each car $i$ parks on its “first attempt”. In other words, if a car encounters either position $n+1$ or position 0 before position $i$, then the entire parking process was unsuccessful. Thus, we have the following theorem giving the expected time for a single car to park, given that all previous cars park.
@@
Let $g^{i}_{s}\coloneqq\mathbb{E}_{s}\left[\tau_{\alpha}^{i}\mid\prod_{j=1}^{i-1}X_{\alpha}^{j}=1\right]$ where we use $\mathbb{E}_{s}$ to denote the expectation of the process starting from $s$. In other words $g^{i}_{s}$ is the expected time it takes for car $i$ to successfully park when starting from position $s$, and thus, $g^{i}_{0}=g^{i}_{i}=0$. Furthermore, letting $\gamma$ be a path from $s$ to $i$, we have that
@@
$\sum_{\ell=1}^{\infty}\sum_{\begin{subarray}{c}\begin{subarray}{c}\begin{subarray}{c}\begin{subarray}{c}|\gamma|=\ell\\ \gamma_{0}=s\end{subarray}\\ \gamma_{1}=s+1\end{subarray}\\ \vdots\end{subarray}\\ \gamma_{\ell}=i\end{subarray}}^{\infty}\left((\ell-1)\Pr\left[\gamma\cap\left\{Z_{\tau_{\alpha}^{i}}^{i}=i,Z_{0}^{i}=s\right\}\right]+\Pr\left[\gamma\cap\left\{Z_{\tau_{\alpha}^{i}}^{i}=i,Z_{0}^{i}=s\right\}\right]\right)$
$=\sum_{\ell=1}^{\infty}p\sum_{\begin{subarray}{c}\begin{subarray}{c}\begin{subarray}{c}|\beta|=\ell-1\\ \beta=s+1\end{subarray}\\ \vdots\end{subarray}\\ \beta_{\ell}=i\end{subarray}}^{\infty}\left((\ell-1)\Pr\left[\beta\cap\left\{Z_{\tau_{\alpha}^{i}}^{i}=i,Z_{0}^{i}=s+1\right\}\right]+\Pr\left[\beta\cap\left\{Z_{\tau_{\alpha}^{i}}^{i}=i,Z_{0}^{i}=s+1\right\}\right]\right)$
$=\sum_{\ell=1}^{\infty}p\sum_{\begin{subarray}{c}\begin{subarray}{c}\begin{subarray}{c}|\beta|=\ell-1\\ \beta=s+1\end{subarray}\\ \vdots\end{subarray}\\ \beta_{\ell}=i\end{subarray}}^{\infty}(\ell-1)\Pr\left[\beta\cap\left\{Z_{\tau_{\alpha}^{i}}^{i}=i,Z_{0}^{i}=s+1\right\}\right]+\sum_{\begin{subarray}{c}\begin{subarray}{c}\begin{subarray}{c}|\beta|=\ell-1\\ \beta=s+1\end{subarray}\\ \vdots\end{subarray}\\ \beta_{\ell}=i\end{subarray}}^{\infty}\Pr\left[\beta\cap\left\{Z_{\tau_{\alpha}^{i}}^{i}=i,Z_{0}^{i}=s+1\right\}\right]$
By the definition of $w^{i}_{s}$ and $g^{i}_{s}$ this last line is just
$pw^{i}_{s+1}\left(g_{s+1}^{i}+1\right)\,.$
@@
The set of parking functions that are weakly increasing are in bijection with those that are weakly decreasing via the involution that takes $i\mapsto n-i+1$, with the added condition that we reverse the order of entrance into the street. That is, the cars would now enter from the right, and perform the parking protocol moving toward spot 1. This involution and change of parking direction in the parking protocol also give a bijection between the set of parking functions whose outcome is the identity permutation and the set of (right to left) parking functions whose outcome is the longest word, namely the permutation $n(n-1)(n-2)\cdots 321$ written in one-line notation. We denote the set of all such parking functions by $\mathsf{PF}_{n}(n\cdots 321)$ and note that they are characterized as the set of parking functions $\beta\in[n]^{n}$ such that the $i$th entry $\beta_{i}$ is at least ni+1n-i+1 for all $i\in[n]$. By way of this bijection, all of our previous arguments and results for the probabilistic parking protocols (both in the open boundary and the unbounded models) work analogously when starting with the set $\mathsf{PF}_{n}(n\cdots 321)$, of parking functions with outcome $n\cdots 321$.
@@ 4.2. Combinatorial Derivation of Expected Time to Park
First, we consider the number of paths from spot $1$ to spot $i$ in $t$ steps. Following [zigzag], we let $a_{j,k}$ be the number of paths to spot $k+1$ using $j$ steps to the left. We can think of $k$ as the net number of spots that the car has moved to the right, so $k+2j$ is the total number of steps the paths enumerated by $a_{j,k}$ requires. There is one path to any spot using only right steps, so $a_{0,k}=1$. We also have that, analogously to (3.1), the number of paths can be determined by the recurrence relation $a_{j,k}=a_{j,k-1}+a_{j-1,k+1}$. This is because a path to spot $s$ can be obtained from either a path landing at spot s1s-1 followed by one right step, or a path landing in spot $s+1$ followed by one left step. Unlike in (3.1) however, a path in the open boundary case cannot end in spot 0. Lastly, since car $i$ always parks in spot $i$ before reaching spot $i+1$, we have that the number of paths landing in spot $k+1$ in $k+2j$ steps is the sequence $a_{j,k}$ from [zigzag] given by
[…]
For any $i$, László Németh and László Szalay refer to the sequence $(a_{j,k})_{j=0}^{\infty}$ as $A_{k}^{(i)}$. Using this notation, the number of paths that car $i$ has starting at spot 1, moving between spot $1$ and i1i-1 in the first k+2j1k+2j-1 steps of the walk and ending at spot $i$ at step $k+2j$ step is the $j$th element in the sequence $A_{i+2}^{(i+2)}$. We use this to find the expected time it takes for car $i$ to travel from spot 1 to spot $i$, given that car $i$ eventually does park which we again denote by $g_{1}^{i}$. We remark that the approach in this section is very similar to that used by McMullen [mcmullen2020drunken], however in their case the terms of the sequence $A_{k}^{(i)}$ are Catalan convolutions because of the different boundary condition. Using the recurrence relation in (4.4), we have the following theorem.
@@
We have remarked that the $j$th element of $A_{i-3}^{(i-3)}$ is the number of paths from 1 to $i$ with $j$ steps to the left. Since we know any path from $1$ to $i$ has i1i-1 more right steps than left steps, we have j+j+i1j+j+i-1 total steps in the walk. The probability of such a path occurring from the set of any path is $q^{j}$ times $p^{j+i-1}$ for the left and right steps, respectively. Where we again use $w_{1}^{i}$ to denote the probability of car $i$ parking in spot $i$ before leaving the segment. Thus, since we only want to sample from the functions that park we multiply $p^{k+i-1}q^{k}$ by $1/w_{1}^{i}$ for the probability of selecting a path with $j$ left steps and 2j+i12j+i-1 total steps, from those which never go to the left of spot 1. ∎
@@
Given a parking function $\alpha=(\alpha_{1},\alpha_{2},\ldots,\alpha_{n})\in\mathrm{PF}_{n}(\mathrm{id})$, the expected time for car $i$ to park in spot $i$ starting from spot i1i-1 provided it parks and all the previous cars have parked under the probabilistic parking protocol with open boundaries is given by
@@ 5. The Parking Process is Negatively Correlated
In this section, we begin by proving that the indicator random variables $X^{i}_{\alpha}$, which are equal to one if car $i$ parks during the probabilistic parking protocol with open boundaries starting with preference list $\alpha$, and zero otherwise, are pairwise negatively correlated.
@@
In the case of our parking protocol with open boundaries, this is exactly as one might expect. Once a car parks, there are fewer spots in the parking lot and so the probability a car is able to park before leaving the segment should only remain the same or decrease. Note that in the context of the unbounded model on the other hand, the indicator random variables (defined analogously to $X^{i}_{\alpha}$) are not negatively correlated. This is because, a car parks under the unbounded model starting with $\alpha\in\mathrm{PF}_{n}(\mathrm{id})$ if and only if it is able to hit position $i$ before drifting off to infinity. However in the case that car $i$ drifts off to infinity, none of the cars in the set $[i+1,n]$ are able to begin the parking process, and thus park with probability zero. We will also give a proof of this fact at the end of the section.
@@
Consider two indices $i,j\in[n]$ with $i<j$. First notice that, while the probability of car $i$ parking is not influenced by car $j$ whenever $i<j$, the events are still correlated. For the left hand side of (5.1), we have
$\Pr[X^{i}_{\alpha}X^{j}_{\alpha}=1]=\Pr[X^{i}_{\alpha}=1]\Pr[X^{j}_{\alpha}=1|X^{i}_{\alpha}=1]\,.$
Thus, what is left to show is that
[…]
Let $\operatorname{Out}(i)$ be the random variable denoting the parking spot in which car $i$ parks as a result of the process, or 0 if car $i$ does not manage to park. Notice that
[…]
where for each $k$,
$\Pr\left[X_{j}^{\alpha}=1\mid\operatorname{Out}(i)=k\right]\leq\Pr\left[X_{j}^{\alpha}=1,\operatorname{Out}(j)\neq k\right]\leq\Pr\left[X_{j}^{\alpha}=1\right]\,.$
Thus, we have
$\begin{split}\Pr\left[X_{j}^{\alpha}=1\mid X_{i}^{\alpha}=1\right]&\leq\Pr\left[X_{j}^{\alpha}=1\right]\sum_{k=1}^{n}\Pr\left[\operatorname{Out}(i)=k\mid X_{i}^{\alpha}=1\right]\\ &=\Pr\left[X_{j}^{\alpha}=1\right]\end{split}$
this shows that (5.2), is indeed an inequality, and concludes the proof.
@@
(5.3) $\prod\limits_{m=i_{1}}^{i_{s}}\Pr\left[X^{m}_{\alpha}=1\mid\prod\limits_{j=i_{1}}^{m-1}X^{j}_{\alpha}=1\right]\stackrel{{\scriptstyle?}}{{\leq}}\prod\limits_{i\in S}\Pr[X^{i}_{\alpha}=1]$
as the left hand side of the above inequality is precisely equal to $\Pr\left[\prod\limits_{i\in S}X^{i}_{\alpha}=1\right]$ by the definition of conditional probability and the probabilistic parking protocol. Thus, we proceed by induction on the size of the set $S$. For the base case, when $S$ is of size 1, the desired inequality holds directly. In addition, for $S$ containing two cars, Theorem 5.3 proves the inequality holds. Now assume for induction that the result holds for $|S|=k$. We now show the inequality holds for $|S|=k+1$. Again, let $\bar{S}\coloneqq\{i_{1},\dots,i_{k+1}\}$ where $i_{1}<\dots<i_{k+1}$. Thus, using the induction hypothesis, we have
[…]
and what is left to show is that
[…]
Notice that we can partition the state space, based only on which spots are occupied when car $j$ attempts to park. Thus, let $O_{K}$ denote the event that the spots in the set $K\subseteq[n]$ are occupied. We have
$\Pr\left[X_{i_{k+1}}^{\alpha}=1\mid\prod_{m=1}^{k}X_{i_{m}}^{\alpha}=1\right]=\sum\limits_{\begin{subarray}{c}K\subseteq[n]:\\ |K|=k\end{subarray}}\Pr\left[X_{i_{k+1}}^{\alpha}=1\mid O_{K}\right]\cdot\Pr\left[O_{K}\mid\prod_{m=1}^{k}X_{i_{m}}^{\alpha}=1\right]\,$,
where one can see that
$\Pr\left[X_{i_{k+1}}^{\alpha}=1\mid O_{K}\right]\leq\operatorname{Pr}\left[X_{i_{k+1}}^{\alpha}=1\right]$
holds. Thus we have,
$\begin{gathered}\operatorname{Pr}\left[X_{i_{k+1}}^{\alpha}=1\mid\prod_{m=1}^{k}X_{i_{m}}^{\alpha}=1\right]\leq\operatorname{Pr}\left[X_{i_{k+1}}^{\alpha}=1\right]\sum\limits_{\begin{subarray}{c}K\subseteq[n]:\\ |K|=k\end{subarray}}\operatorname{Pr}\left[O_{K}\mid\prod_{m=1}^{k}X_{i_{m}}^{\alpha}=1\right]\\ =\operatorname{Pr}\left[X_{i_{k+1}}^{\alpha}=1\right]\end{gathered}$
which yields the desired result. ∎
@@
Furthermore, by Theorem 3.1 we know that the probability that any particular car (which is not the first car) parks with probability 1, depends on the parameter $p$. Thus, it is not true that this inequality always holds, and so the probabilistic parking process with the unbounded boundary condition, is not negatively correlated in general. ∎